# Underwater SONAR Image Classification and Analysis using LIME-based Explainable Artificial Intelligence

### v1

@@ VI-B2 Balanced Data
The process of image classification using transfer learning is performed using this balanced data using stratified sampling. The accuracy achieved by each model during the training process with balanced data on the test dataset and validation dataset is summarised in Table XIII and depicted in the Fig. 9. These results highlight the performance of DenseNet121, which achieved an impressive accuracy of 98.21%, DenseNet201 - 97.55%, NASNetLarge - 97.10%, MobileNetV2 and InceptionResNetV2 - 96.65%, Xception - 95.76%, InceptionV3 and NASNetMobile - 95.66%. VGG19 - 94.87%, VGG16 - 94.87 and ResNet50 - 84.40%, all surpassing the 95% accuracy mark.
TABLE XIII: Accuracy of Deep Learning Models developed utilizing stratified sampling approaches with balanced data.
@@
The classification report, as shown in Table XIV and the Fig. 10(b), indicates that the class-wise accuracy has improved. By using a stratified approach on balanced data, the accuracy for “human”, “plane” and “ship” has reached 100%, 98% and 97% respectively. This represents an improvement from the previous model’s accuracy developed using balanced data through random sampling techniques for “plane” (from 93% to 98%) and for “ship” (from 94% to 97%).
TABLE XIV: Classification Report for balanced Model using stratified sampling approach-DenseNet121.
Class Precision Recall F1-Score Accuracy
Fish 1.00 1.00 1.00 1.00
Human 1.00 1.00 1.00 1.00
Mine 1.00 1.00 1.00 1.00
Plane 0.99 0.97 0.98 0.98
Seafloor 1.00 0.99 0.99 1.00
Ship 0.97 1.00 0.99 0.97
(a) DenseNet121 - Unbalanced Model
(b) DenseNet121 - Balanced Model
These insights emphasize the critical role of choosing the correct sampling approach and balanced data along with well-suited, pre-trained models. Among the options considered, DenseNet121 from the stratified approach using balanced data stood out as the most fitting choice for our classification task, demonstrating outstanding validation accuracy of 99.33% and a test accuracy of 98.21%, hence it is considered the de-facto model for our following XAI study.
VI-C Explaining the predictions from the classifier using SP-LIME
As a result of various combinations of random vs. stratified sampling strategies as well as balanced and unbalanced approaches (as mentioned in previous Sections VI-A and VI-B), we developed altogether forty-four distinct classification models, with forty of them having an accuracy of more than 95%. We chose the model DenseNet121 with the highest accuracy and superior generalisation abilities for each class developed using stratified approach on balanced data and incorporated the post-hoc SP-LIME technique, a variation of LIME, to enhance explainability.
Although LIME can provide valuable insights, it is important to carefully choose which instances perturb, since perturbing all the pixels in the image demands more time and computation, hence selected instances are perturbed such as super-pixels, viz. SP-LIME. This is because users may not have the time or capacity to review a large number of explanations, as explained in Section IV-D.
@@ VI-C Explaining the predictions from the classifier using SP-LIME
Figure 12: Explaining the prediction made by DenseNet121 neural network (Balanced Model) using SP-LIME. The top 3 classes predicted are ”Plane” (p = 0.100), ”Ship” (p = 0.000) and ”Human” (p = 0.000)
@@ VI-D Ablation Study
VI-D1 Explainer with balanced and unbalanced models
In this experiment, the impact of the SP-LIME explainer with sampling and data balancing strategies is analysed. Both the random and stratified sampling along with balanced vs unbalanced datasets using our top-performing model are studied. Table XV summarises the visual explanation along with the original image.
TABLE XV: SP-LIME with different models: First row: sampling techniques, second row: model’s balance, third-row predicted class: plane, and fourth-row predicted class: ship, and the corresponding visual explanation images from that particular class.
Table XV, it can be observed that the balanced model from the stratified sampling approach provides the best visual explanation, with a focus on the area with the exact mask. Compared to the random sampling technique, stratified sampling outperformed in selecting the target object precisely. Similarly, it is seen that the unbalanced model from both approaches results in more shadow regions, compared to the balanced model. For the plane and ship prediction, it is clear that the mask from the stratified balanced model outperforms the other models in explaining the tail&wing and the relevant ship regions, respectively.
VI-D2 Hyper-parameter of the Explainer- Number of Features
This section tries to understand the impact of number of features as a hyperparameter towards LIME explanation. Since SP-LIME perturbs at the super-pixel level while LIME perturbs each pixel, we assess the computational time taken for each approach, to better understand the performances and computational requirements of LIME and SP-LIME. We test LIME and SP-LIME’s performance in a number of different parameter settings, specifically investigating the number of features, tweaking to 10, 25, 50, 100, 150, 200, 250, 300, 400, and 500. Although SP-LIME has its own hyper-parameters, we have considered the following parameters as fixed as follows: super-pixel algorithm: quickshift, kernel size: 2, max distance: 100, and ratio: 0.1.
The number of features denote the quantity of features or input variables that the explainer will perturb in order to produce a local explanation. The absence of a large class group, coupled with relatively low variability in the dataset, diminishes the computational expense sensitivity to changes in the number of features for both LIME and SP-LIME. Despite this, the choice of the feature count significantly influences the effectiveness of the explanations provided for the classification decisions (refer to Fig. 13) and Table XVI.
TABLE XVI: Comparison with 100 Samples on LIME & SP-LIME
@@ VI-D2 Hyper-parameter of the Explainer- Number of Features
From Table XVI, it can be observed that if more features are considered for both LIME and SP-LIME, then it may result in more computational time requirement spent calculating the mask. The comparative visualization analysis of LIME and SP-LIME by varying the number of features is given in Fig. 13. It can be observed that Fig. 13 (a) from LIME and Fig. 13 (b) from SP-LIME with 10 features outperform the rest (c) and (d) with 500 features. Further, SP-LIME takes less time to compute but gives a comparable explanation to LIME, which can be further tuned with some changes in parameters.
VI-D3 Hyper-parameter of the Explainer- Number of Samples
The number of samples represents the number of synthetic samples generated by the explainer for each perturbation. We fix the number of features to 10, vary the number of samples with values ranging from 100 to 1000, and analyse the performance of our XAI model. The computational time associated with each combination of parameters are shown in Table XVII.
@@ VI-D3 Hyper-parameter of the Explainer- Number of Samples
TABLE XVII: Comparison with 10 Features on LIME & SP-LIME.
@@
From Table XVII, it can be observed that the number of samples also influences the explainer, which results in an increase in computation with an increase in the number of samples. From Fig. 14, it can be noted that the explanation from SP-LIME is better than LIME since the LIME captured more regions, which is of no interest with more computational resources and more time, but on the other hand, the SP-LIME masked the zone, which is very close to the region of interest. Depending on the dataset and the computational constraints available, a trade-off between the number of features and samples should be made for a better explanation. So we chose 10 features and 500 samples and applied them to both LIME and SP-LIME. The LIME explainer took 31.61 seconds. On the other hand, the SP-LIME explainer took 20 seconds and produced comparable results (refer to Fig. 15).
@@
VI-D4 SP-LIME and the corresponding Super-pixel Algorithm
@@
VI-D5 Impact of explainer on different classes
@@
The explanation mask accuracy differs due to training data. The original “human” dataset had 34 images, and augmentation added 498 images to this dataset. However, this augmentation may have caused the images to differ, causing the model to learn patterns rather than accurately capture human objects’ intrinsic shapes. Hence, the model’s performance suffers, and the explainer struggles to provide clear explanations. For the ‘fish’ and ‘mine’ classes, the dataset is of poor quality with data limitations and similarity, so the model failed to provide meaningful explanations. On the other hand, for the classes plane, ship, and seafloor (refer Fig. 18(a), Fig. 18(b), and Fig. 18(c)), promising explanations are achieved due of the quality and quantity of images used for training. We envisage that image enhancement techniques such as SRGAN [86] and attention techniques could be employed to enhance image quality and to direct explainer towards the area of interest.
@@
VI-E State of the Art Comparison
To showcase the efficacy of our classification model, a thorough comparison was conducted with existing models, evaluating their respective accuracies. In underwater SONAR imagery, there are very few research works as follows, Divya et al. [87] employed ResNet50 and ensemble techniques on seabed KLSG data, achieving a notable accuracy of 90%. In a study by Najib et al. [88] on NAT III, a DCNN yielded an impressive accuracy of 98.44%. Wang et al. [89], utilizing AWCNN, achieved an accuracy of 86.90%. Nguyen et al. [33], working with CKI and TDI datasets, utilized GoogLeNet and attained an accuracy of 91.60%. Chungath et al. [8], combining seabed and custom datasets and employing ResNet34, reported an accuracy of 97.33%. Fernandes et al. [90] utilized a tailor-made CNN on a custom dataset, achieving a maximum accuracy of 99%. Luo et al. [7], using a shallow CNN with seabed sediment data, obtained an accuracy of 94.40%. Zhou et al. [91] employed SCTD and achieved an accuracy of 95.84 through IMDNet. The results, as illustrated in Table XVIII, unequivocally highlight the superior performance of our customised DenseNet121 model, positioning it at the forefront of the evaluated models with an accuracy of 98.21%.
TABLE XVIII: State of the Art (SOTA) Comparison - Image Classifier.
@@ VI-E State of the Art Comparison
TABLE XIX: State of the Art (SOTA) Comparison: Image Explainer.
@@
For the comparison of our explainer in our research paper, we reviewed three key papers that contribute to the field of explainable AI for synthetic aperture imagery, tabulated in Table XIX. In the field of SONAR imagery, Walker et al. [55] used LIME to learn about important features that play a part in improving performance and to learn how important it is to balance data for fine-tuning deep learning models in classifying SAS seafloor images. In 2023, Oveis et al. [92] conducted a study using LIME, for target recognition with synthetic radar images. Additionally, Pannu et al. [40] also utilized LIME for a similar purpose. Utilizing radar images, Pannu et al.’s (2020) study used LIME to perform target classification, Both studies utilized the MSTAR dataset. Since there are no specific metrics for explaining the model, the usefulness of these explainer models depends entirely on the user’s perspective. We utilized LIME to provide explanations for the predictions made by an underwater SONAR image classifier, thereby improving the trust and effectiveness of the black box model.
VII Conclusion and Future Works

### v2

@@ 6.2.2 Balanced Data
The process of image classification using transfer learning is performed using this balanced data using stratified sampling. The accuracy achieved by each model during the training process with balanced data on the test dataset and validation dataset is summarised in Table 14 and depicted in the Fig. 10. These results highlight the performance of DenseNet121, which achieved an impressive accuracy of 98.21%, DenseNet201 - 97.55%, NASNetLarge - 97.10%, MobileNetV2 and InceptionResNetV2 - 96.65%, Xception - 95.76%, InceptionV3 and NASNetMobile - 95.66%. VGG19 - 94.87%, VGG16 - 94.87 and ResNet50 - 84.40%, all surpassing the 95% accuracy mark.
Table 13: Classification Report for balanced Model using stratified sampling approach-DenseNet121.
Class Precision Recall F1-Score Accuracy
Fish 1.00 1.00 1.00 1.00
Human 1.00 1.00 1.00 1.00
Mine 1.00 1.00 1.00 1.00
Plane 0.99 0.97 0.98 0.98
Seafloor 1.00 0.99 0.99 1.00
Ship 0.97 1.00 0.99 0.97
Table 14: Accuracy of Deep Learning Models developed utilizing stratified sampling approaches with balanced data.
@@
The classification report, as shown in Table 13 and the Fig. 9(b), indicates that the class-wise accuracy has improved. By using a stratified approach on balanced data, the accuracy for “human", “plane" and “ship" has reached 100%, 98% and 97% respectively. This represents an improvement from the previous model’s accuracy developed using balanced data through random sampling techniques for “plane" (from 93% to 98%) and for “ship" (from 94% to 97%). These insights emphasize the critical role of choosing the correct sampling approach and balanced data along with well-suited, pre-trained models. Among the options considered, DenseNet121 from the stratified approach using balanced data stood out as the most fitting choice for our classification task, demonstrating outstanding validation accuracy of 99.33% and a test accuracy of 98.21%, hence it is considered the de-facto model for our following XAI study.
6.3 Explaining the predictions from the classifier using SP-LIME
As a result of various combinations of random vs. stratified sampling strategies as well as balanced and unbalanced approaches (as mentioned in previous Sections 6.1 and 6.2), we developed altogether forty-four distinct classification models, with forty of them having an accuracy of more than 95%. We chose the model DenseNet121 with the highest accuracy and superior generalisation abilities for each class developed using stratified approach on balanced data and incorporated the post-hoc SP-LIME technique, a variation of LIME, to enhance explainability.
Although LIME can provide valuable insights, it is important to carefully choose which instances perturb, since perturbing all the pixels in the image demands more time and computation, hence selected instances are perturbed such as super-pixels, viz. SP-LIME. This is because users may not have the time or capacity to review a large number of explanations, as explained in Section 4.4.
@@ 6.3 Explaining the predictions from the classifier using SP-LIME
Figure 12: Explaining the prediction made by DenseNet121 neural network (Balanced Model) using SP-LIME. The top 3 classes predicted are "Plane" (p = 0.100), "Ship" (p = 0.000) and "Human" (p = 0.000)
@@ 6.4 Ablation Study
6.4.1 Explainer with balanced and unbalanced models
In this experiment, the impact of the SP-LIME explainer with sampling and data balancing strategies is analysed. Both the random and stratified sampling along with balanced vs unbalanced datasets using our top-performing model are studied. Table 15 summarises the visual explanation along with the original image.
Table 15: SP-LIME with different models: First row: sampling techniques, second row: model’s balance, third-row predicted class: plane, and fourth-row predicted class: ship, and the corresponding visual explanation images from that particular class.
Table 15, it can be observed that the balanced model from the stratified sampling approach provides the best visual explanation, with a focus on the area with the exact mask. Compared to the random sampling technique, stratified sampling outperformed in selecting the target object precisely. Similarly, it is seen that the unbalanced model from both approaches results in more shadow regions, compared to the balanced model. For the plane and ship prediction, it is clear that the mask from the stratified balanced model outperforms the other models in explaining the tail&wing and the relevant ship regions, respectively.
6.4.2 Hyper-parameter of the Explainer: Number of Features
This section tries to understand the impact of number of features as a hyperparameter towards LIME explanation. Since SP-LIME perturbs at the super-pixel level while LIME perturbs each pixel, we assess the computational time taken for each approach, to better understand the performances and computational requirements of LIME and SP-LIME.
Table 16: Comparison with 100 Samples on LIME & SP-LIME
@@ 6.4.2 Hyper-parameter of the Explainer: Number of Features
We test LIME and SP-LIME’s performance in a number of different parameter settings, specifically investigating the number of features, tweaking to 10, 25, 50, 100, 150, 200, 250, 300, 400, and 500. Although SP-LIME has its own hyper-parameters, we have considered the following parameters as fixed as follows: super-pixel algorithm: quickshift, kernel size: 2, max distance: 100, and ratio: 0.1. The number of features denote the quantity of features or input variables that the explainer will perturb in order to produce a local explanation. The absence of a large class group, coupled with relatively low variability in the dataset, diminishes the computational expense sensitivity to changes in the number of features for both LIME and SP-LIME. Despite this, the choice of the feature count significantly influences the effectiveness of the explanations provided for the classification decisions (refer to Fig. 13) and Table 16.
From Table 16, it can be observed that if more features are considered for both LIME and SP-LIME, then it may result in more computational time requirement spent calculating the mask. The comparative visualization analysis of LIME and SP-LIME by varying the number of features is given in Fig. 13. It can be observed that Fig. 13 (a) from LIME and Fig. 13 (b) from SP-LIME with 10 features outperform the rest (c) and (d) with 500 features. Further, SP-LIME takes less time to compute but gives a comparable explanation to LIME, which can be further tuned with some changes in parameters.
6.4.3 Hyper-parameter of the Explainer: Number of Samples
The number of samples represents the number of synthetic samples generated by the explainer for each perturbation. We fix the number of features to 10, vary the number of samples with values ranging from 100 to 1000, and analyse the performance of our XAI model. The computational time associated with each combination of parameters are shown in Table 17.
@@ 6.4.3 Hyper-parameter of the Explainer: Number of Samples
Table 17: Comparison with 10 Features on LIME & SP-LIME.
@@
From Table 17, it can be observed that the number of samples also influences the explainer, which results in an increase in computation with an increase in the number of samples. From Fig. 14, it can be noted that the explanation from SP-LIME is better than LIME since the LIME captured more regions, which is of no interest with more computational resources and more time, but on the other hand, the SP-LIME masked the zone, which is very close to the region of interest. Depending on the dataset and the computational constraints available, a trade-off between the number of features and samples should be made for a better explanation. So we chose 10 features and 500 samples and applied them to both LIME and SP-LIME. The LIME explainer took 31.61 seconds. On the other hand, the SP-LIME explainer took 20 seconds and produced comparable results (refer to Fig. 15).
@@
6.4.4 SP-LIME and the corresponding Super-pixel Algorithm
@@
6.4.5 Impact of explainer on different classes
@@
The explanation mask accuracy differs due to training data. The original “human" dataset had 34 images, and augmentation added 498 images to this dataset. However, this augmentation may have caused the images to differ, causing the model to learn patterns rather than accurately capture human objects’ intrinsic shapes. Hence, the model’s performance suffers, and the explainer struggles to provide clear explanations. For the ‘fish’ and ‘mine’ classes, the dataset is of poor quality with data limitations and similarity, so the model failed to provide meaningful explanations. On the other hand, for the classes plane, ship, and seafloor (refer Fig. 18(a), Fig. 18(b), and Fig. 18(c)), promising explanations are achieved due of the quality and quantity of images used for training. We envisage that image enhancement techniques such as SRGAN [81] and attention techniques could be employed to enhance image quality and to direct explainer towards the area of interest.
@@
6.5 State of the Art Comparison
To showcase the efficacy of our classification model, a thorough comparison was conducted with existing models, evaluating their respective accuracies. In underwater SONAR imagery, there are very few research works as follows, Divya et al. [15] employed ResNet50 and ensemble techniques on seabed KLSG data, achieving a notable accuracy of 90%. In a study by Najib et al. [48] on NAT III, a DCNN yielded an impressive accuracy of 98.44%. Wang et al. [80], utilizing AWCNN, achieved an accuracy of 86.90%. Nguyen et al. [53], working with CKI and TDI datasets, utilized GoogLeNet and attained an accuracy of 91.60%. Chungath et al. [10], combining seabed and custom datasets and employing ResNet34, reported an accuracy of 97.33%. Fernandes et al. [20] utilized a tailor-made CNN on a custom dataset, achieving a maximum accuracy of 99%. Luo et al. [41], using a shallow CNN with seabed sediment data, obtained an accuracy of 94.40%. Zhou et al. [87] employed SCTD and achieved an accuracy of 95.84 through IMDNet. The results, as illustrated in Table 18, unequivocally highlight the superior performance of our customised DenseNet121 model, positioning it at the forefront of the evaluated models with an accuracy of 98.21%.
Table 18: State of the Art (SOTA) Comparison - Image Classifier.
@@ 6.5 State of the Art Comparison
Table 19: State of the Art (SOTA) Comparison: Explainer (XAI).
@@
For the comparison of our explainer in our research paper, we reviewed three key papers that contribute to the field of explainable AI for synthetic aperture imagery, tabulated in Table 19. In the field of SONAR imagery, Walker et al. [79] used LIME to learn about important features that play a part in improving performance and to learn how important it is to balance data for fine-tuning deep learning models in classifying SAS seafloor images. In 2023, Oveis et al. [56] conducted a study using LIME, for target recognition with synthetic radar images. Additionally, Pannu et al. [57] also utilized LIME for a similar purpose. Utilizing radar images, Pannu et al.’s (2020) study used LIME to perform target classification, Both studies utilized the MSTAR dataset. Since there are no specific metrics for explaining the model, the usefulness of these explainer models depends entirely on the user’s perspective. We utilized LIME to provide explanations for the predictions made by an underwater SONAR image classifier, thereby improving the trust and effectiveness of the black box model.
7 Conclusion and Future Works